OpenStudy (anonymous):

What is the steady state power dissipation for a driven damped harmonic oscillator with sinusoidal input? Where does the 2 pi factor come from in the equation \[ Q = 2\pi \frac{Energy stored}{Energy dissipated per cycle}\]

OpenStudy (anonymous):

\[Q = \frac{m\omega_0}{b}\] Where b is the damping constant, such that \[F_{damping} = -bv\] The problem I am having is \[Power = Fv = -bv^2\] Where \[v = -\omega A \sin(\omega t + \theta)\] \[E = \int\limits_{0}^{2\pi/\omega} Power(t)dt= -b\omega^2A^2\int\limits_{0}^{2\pi/\omega}\sin^2(\omega t + \theta)dt = -b\omega^2A^2/2\] There's no 2 pi there, and my book indicates there should be a factor of 2pi, and the power of omega should be one, not two. However, there is no period in the book's expression, so substituting 2pi/T for one omega won't do it.

OpenStudy (unklerhaukus):

\[\frac{2\pi}{cycle}=1\]

OpenStudy (anonymous):

I confused the time integral of sin^2 with its average value. The time integral over a period should be\[\pi/\omega\]